Consider all samples of size of 5 from this population:

2, 5, 7, 9, 10, 11, 12


Compute the population mean

Compute the population variance

Compute the population standard deviation

Compute the mean of the sampling distribution of the sample means

Compute the variance of the sampling distribution of the sample means.

Compute the standard deviation of the sampling distribution of the sample means.


1
Expert's answer
2022-01-31T12:05:46-0500

a.

The population mean

"\mu = \frac{2+5+7+9+10+11+12}{7}= 8"


b.

The population variance

"s^2 = \frac{1}{n-1} \sum (x_i - \mu)^2 \\

s^2 = \frac{1}{7-1}((2-8)^2 + (5-8)^2 + (7-8)^2 + (9-8)^2 + (10 -8)^2 + (11 -8)^ + (12 -8)^2) \\

= \frac{1}{6}( 36 + 9 + 1 + 1+ 4 + 9 + 16 ) \\

= 12.66"


c.

The population standard deviation

"s = \sqrt{24.61} = 3.55"


d.

The number of all possible sample = 7C5=21 and the corresponding sample mean = sum of each five values/5



The mean of the sampling distribution of the sample means = 8

The mean of the sampling distribution of the sample means is equal to compare it the mean of the population.


e.

The variance of the sampling distribution of the sample means

"s^2 = \frac{1}{21-1} \times 15.2 = 0.76"


f.

The standard deviation of the sampling distribution of the sample means

"s = \sqrt{0.76} = 0.87"